Hello,
If someone could help me solve for this problem from, the Book: Intermediate Algebra by Mark Dugopolski, I would greatly appreciate it.
(The 2 next to the 4 is squared)
Evaluate expression.
8 - 3 |5 - 4^2 + 1|
Thank you very much

Answer by CubeyThePenguin(3113)

8 - 3|5 - 4^2 + 1|
= 8 - 3|5 - 16 + 1|
= 8 - 3|-10|
= 8 - 3(10)
= 8 - 30
= -22